OpenStudy (anonymous):

-6(2p+4)-30=2(5p+6) how do you do this step by step

OpenStudy (anonymous):

solve for p presumably: distribute -12p-24-30=10p+12 bring p to 1 side, numbers to the other -12p-10p-54 = 12 -22p = 66 p = -3 Questions?

OpenStudy (anonymous):

no an thanks